OpenStudy (anonymous):

i need help with this calculus problem. 1/(x-7) x=-2 x=0 find average rate of change for this function

myininaya (myininaya):

\[\frac{\frac{1}{-2-7}-\frac{1}{0-7}}{-2-0}\]

OpenStudy (anonymous):

u just replace the given x's to the formula?

myininaya (myininaya):

\[Average rate of change=\frac{f(x_1)-f(x_2)}{x_1-x_2}\]

myininaya (myininaya):

so yes